Recent Log Entries

QE Measurement

Log Number: 4104938Logbook(s): Body: *New Spot* Spot Location: (10700.0,18500.0)

Laser Wavelength Atten Setting Laser Power Current HV PS Current QE PV Updated? Update Time/Reason Not Updated (nm) (W) (uA)

Text alert added to 10Amp trim alarms

Log Number: 4104930Logbook(s): Body: 

Received a request to add a text alert if a 10Amp trim goes into MAJOR alarm.
Alerts will be sent to Jim Colman and Kevin Banks.
Restarted the alarm handlers at with the permission of the crew.

Autolog save

Log Number: 4104925Logbook(s): Tags: Body: saver ran from opsl06

Saved: ywang 01-04-23 16:37:50 {A:3 B:5 C:5 E:1047 A=Tune B=Pulse C=Pulse D=BS 0R08=0.00 0R08 dump in. dump(s) in.

Separator Update

Log Number: 4104918Logbook(s): Entry Makers: wissmannxgomezBody: 

Pass 4 and Pass 5 Horizontal and Vertical separators were started and operated at full power to verify their operation and optimize tuning. All separators were then tuned off but are ready
for an anticipated easy restart when they are needed,

iocnl6mag restored

Log Number: 4104898Logbook(s): Tags: Body: restorelib 2.5 Target restore date was 2023-01-04 16:06:26 The output from the myRestore execution was as follows: iocnl6mag magnet_ioc magnet_MAG@2023-01-04 16:06 182(182) restored Pause: 2 second(s) magnet_HVAL@2023-01-04 16:06 182(182) restored Pause: 2 second(s) magnet_LOOP@2023-01-04 16:06 182(182) restored Pause: 2 second(s) magnet_CMPQ@2023-01-04 16:06 380(380) restored magnet_TrimRack@2023-01-04 16:06 12(12) restored Procedure return status: Success system@2023-01-04 16:06 2(2) restored E

iocwa1mag restored

Log Number: 4104897Logbook(s): Tags: Body: restorelib 2.5 Target restore date was 2023-01-04 16:06:53 The output from the myRestore execution was as follows: iocwa1mag magnet_ioc magnet_MAG@2023-01-04 16:06 147(147) restored Pause: 2 second(s) magnet_HVAL@2023-01-04 16:06 135(135) restored Pause: 2 second(s) magnet_LOOP@2023-01-04 16:06 135(135) restored Pause: 2 second(s) magnet_CMPQ@2023-01-04 16:06 274(274) restored magnet_TrimRack@2023-01-04 16:06 8(8) restored Procedure return status: Success wa1magFlow@2023-01-04 16:06 4(4) restored

Follow-up Re: Lambertson Viewer

Log Number: 4104885Logbook(s): Body: 

Matt Mueller and I reconfigured the shielding around the Lambertson viewer camera per Joe Gubeli's shielding design using 30% borated ploy and lead. We also installed a shadow shield upstream of the Lambertson magnet to help prolong the camera life. Dose rods were also installed around the camera to track does rates.

NLinac 1L12 Cooldown Complete

Log Number: 4104883Logbook(s): Entry Makers: weaksmccryoBody: 

1L12 Cooldown complete.

Currently filling to 20%, will hold at 20% overnight to allow reconfiguration for 2K pumpdown in the AM.

COMMENT ON updated field control firmware for 1L22

Log Number: 4104882Logbook(s): Body: 

we are not planning to make any software changes. That was the reason SW was not selected.
software - changes at the EPICS and IOC level
firmware - I should probably name this gateware as we are updating the FPGA code only

Pages